I need to understand how to solve this equation and to find the value of x.
4x+x+(4x-20)+(2x-10)=1,202

Answer by elima(1433)

4x+x+(4x-20)+(2x-10)=1,202
First of all you don't need all the parenthesis. They but them there to confuse you. So rewrite the proplem without them:
4x+x+4x-20+2x-10=1202
Now you need to add like terms:
Gather them together:
4x+x+4x+2x-20-10=1,202
11x-30=1202
Now you add 30 to both sides:
11x-30+30=1202+30
11x=1232
Now divide both sides by 11:
11x/11=1232/11
x=112
Hope you get it =)